### v1

@@ 1 Introduction
Formally, we study a two-sided, dynamic matching market, in which agents on both sides of the economy arrive over time. We assume arrivals are deterministic, that is, agents can perfectly foresee when each agent arrives in the economy. A dynamic matching specifies who matches with who in each period, subject to two constraints: agents cannot be matched before they arrive, and agents can have at most one matching partner on the opposite side. Agents are discounted utility maximizers, and their payoff from a dynamic matching solely depends on who they match with and when. Our goal is to study solution concepts for this economy.
@@
Whereas consistency is defined relative to the candidate matching, an immediate corollary is that the following conjectures always lead to a nonempty solution concept. Suppose that each agent’s conjectures only rule out continuation matchings that do not conform with the solution concept in the continuation. For instance, in a two-period economy, these conjectures specify that in the last period only stable matchings among the remaining unmatched agents and the new arrivals are possible. These conjectures always satisfy consistency and Theorem 4.1 implies they lead to a non-empty solution concept. That is, so long as the agents conjecture that the economy will conform to the solution concept in the continuation, the set of matchings that conform to the solution concept from today onward is nonempty.
@@
Our paper builds on the literature on two-sided matching with externalities (Li, 1993; Sasaki and Toda, 1996; Hafalir, 2008) to contribute to the literature on two-sided dynamic matching markets with irreversible matching (Doval, 2022; Nicolò et al., 2023).
@@
The literature on cooperative solution concepts for dynamic matching markets highlights the role in defining solution concepts of agents’ conjectures about what matching would ensue after a block (see, e.g., Kurino, 2009; Kadam and Kotowski, 2018; Kotowski, 2019; Altinok, 2019; Liu, 2023 for repeated matching markets and Doval, 2022; Nicolò, Salmaso, and Saulle, 2023 for irreversible matching, and Doval (Forthcoming) for a survey treatment of this literature). Because solution concepts in this literature are typically predicated on perfection requirements–i.e., the same solution concept applying at the appropriately defined continuation–agents’ conjectures are defined jointly with the solution concept. We follow a complementary approach in which we take agents’ conjectures as given and use them to define a solution concept. This allows us to identify a property of conjectures–consistency–that guarantees the corresponding solution concept is nonempty, and underlies existence results in the literature (cf. Doval, 2022). We also propose new solution concepts that refine those existing in the literature, illustrating how our result can be used as a recipe of sorts to define stability notions for dynamic matching markets.
Recently, Nicolò, Salmaso, and Saulle (2024) study a third refinement of dynamic stability (Doval, 2022). Like the conjectures in dynamic stability, the conjectures in the solution concept in (Nicolò et al., 2024) are built jointly with the solution concept. In particular, the agents rule out all matchings that do not abide by the solution concept from tomorrow onward. Unlike dynamic stability, the solution concept in Nicolò et al. (2024) imposes more structure on what matchings the agents can conjecture will ensue in the current period when considering remaining unmatched: The authors define conjectures in a given period through an iterative procedure which in each stage, rules out current period matchings that do not satisfy some minimal stability requirements given the conjectures at that stage.Footnote 2Footnote 2Footnote 2Whereas the main focus of the paper is on the conjectures that obtain as the limit of this procedure, by stopping the procedure at a given stage the authors can study stability under different levels of sophistication (Hafalir, 2008). It is possible to show the conjectures they construct satisfy consistency. Following Doval (2022), their proof is based on showing that the analogue for their solution concept of the candidate matchings we define in Section 4 satisfy the conditions of their solution concept. Furthermore, the authors show via an example that the rational expectations solution may fail to exist in a two-sided dynamic matching market.
@@ 2 Model
A length-$T$ economy, denoted $E^{T}=(A_{1},B_{1},\dots,A_{T},B_{T})$, consists of two sequences, $\{A_{1},\dots,A_{T}\}$ and $\{B_{1},\dots,B_{T}\}$, of disjoint subsets of $\mathscr{A}$ and $\mathscr{B}$, respectively.Footnote 3Footnote 3Footnote 3It is without loss to assume the sets $(A_{1},\dots,A_{T})$ are disjoint as we do not rule out indifferences. Thus, $a$ denotes a unique individual in the economy, but $a$ may be payoff-equivalent to other arriving agents on side $A$. The same applies for the sets $(B_{1},\dots,B_{T})$ being disjoint. For $t\in\{1,\dots,T\}$, $A_{t}$ are the side $A$-agents that arrive in period $t$, and $B_{t}$ are the side $B$-agents that arrive in period $t$. For any $t\leq T$, let $\overline{A}_{t}=\cup_{s=1}^{t}A_{s}$ denote the implied arrivals on side $A$ through period $t$; similarly, let $\overline{B}_{t}=\cup_{s=1}^{t}B_{s}$ denote the implied arrivals on side $B$ through period $t$.
@@
We close the model by defining the agents’ preferences. Throughout, we assume agents discount the future.Footnote 4Footnote 4Footnote 4Discounted utility simplifies notation but it is not needed for any of our results. In fact, all of our results hold if agents have ordinal preferences over the matchings, so long as the preferences depend only on their own matching outcome and they are dynamically consistent. Formally, each $a\in\mathscr{A}$ defines a discount factor $\delta_{a}\in[0,1]$ and a Bernoulli utility, $u(a,\cdot):\mathscr{B}\cup\{a\}\mapsto\mathbb{R}$. Similarly, each $b\in\mathscr{B}$ defines a discount factor $\delta_{b}\in[0,1]$ and a Bernoulli utility, $v(\cdot,b):\mathscr{A}\cup\{b\}\mapsto\mathbb{R}$. We normalize to 00 the payoffs of remaining single for one period. That is, for all $a\in\mathscr{A}$ and all $b\in\mathscr{B}$, $u(a,a)=v(b,b)=0$.
@@ 3 A solution concept given conjectures
Because the conjectures represent what $k$ expects to obtain should they remain unmatched in period $t$, each conjectured matching $\overline{m}$ satisfies $\overline{m}_{t}(k)=k$. Each conjectured matching $\overline{m}\in\varphi_{t}(m^{t-1},k)$ describes both the matching that ensues in period $t$, $\overline{m}_{t}$, and the continuation matching $(\overline{m}_{t+1},\dots,\overline{m}_{T})$. Whereas agent $k$’s matching outcome is determined by the continuation matching $(\overline{m}_{t+1},\dots,\overline{m}_{T})$, specifying who matches in period $t$ is necessary to define what matchings are feasible from period $t+1$ onward. For this reason, a conjecture describes the matching starting from period $t$ onward. That the outcomes of the other period-$t$ agents matters to define the set of matchings $k$ may face from period $t+1$ onward implies an externality exists in this economy, even if no direct payoff externalities exist.
@@
In words, a matching $m$ is a $\Phi$-solution for $E^{T}$ if in each period the following holds. First, each agent $k$ who is available to match in period $t$ prefers $m$ to their worst conjectured matching (parts 1 and 2 in 6). Note that this implies that if $m_{t}(k)\neq k$, then $k$ prefers their matching partner to waiting to be matched, and if $m_{t}(k)=k$, then $k$ has no objections to the continuation matching, $(m_{s})_{s=t+1}^{T}$.Footnote 5Footnote 5Footnote 5In a static economy, only the first kind of block may exist. The second block only shows up in a dynamic economy (see Doval, 2022 for a discussion). For instance, if the continuation matching is not self-enforcing, an agent may object to it. As we discuss later on, consistency of the conjectures rules is very much related to the second class of blocks. In both cases, we say $k$ prefers to be available to match in period $t$. Second, no pair of agents available to match in period $t$ prefer to match with each other to matching according to $m$ (part 3 in 6). Note that $(a,b)$ being available to match in period $t$ does not imply $a$ and $b$ arrive in the same period, but that under matching $m$ they meet in period $t$ and are able to block $m$.
@@
Whereas 6 is fairly general, it is worth noting that the definition rules out solution concepts in which the conjectures depend on the matching under consideration beyond the induced economy in period $t$. For instance, when considering whether to block matching $m$ in period $t$, if agent $k$ presumes all other contemporaneous agents–except perhaps for $k$’s matching partner–remain matched according to $m_{t}$, then $k$’s conjectures depend on the outcome specified by $m$ in period $t$. (These conjectures are referred to as passive conjectures; see Chowdhury (2004) and Doval (2015).) Instead, 6 presumes we can define the conjectures without reference to any specific matching. The matching under consideration only matters inasmuch as it determines who is available to match in each period, and hence whose conjectures one should check in each period.
@@
    For all $a\in A_{1}$, a matching $\overline{m}\in\varphi_{1}(m^{0},a)$ exists such that $U_{1}(a,m)\geq U_{1}(a,\overline{m})$,
@@
    For all $b\in B_{1}$, a matching $\overline{m}\in\varphi_{1}(m^{0},b)$ exists such that $V_{1}(b,m)\geq V_{1}(b,\overline{m})$,
@@
    $(m_{t})|_{t=2}^{T}\in\Sigma_{T-1}^{\Phi}(E_{2}^{T}(m_{1}))$.Footnote 7Footnote 7Footnote 7We abuse notation slightly. Whereas $\Sigma_{2}^{\Phi}(E_{2}^{T}(m_{1}))$ defines a matching only for the agents who can match from $t=2$ onward–i.e., it is an element of $\pazocal{M}_{T-1}(E_{2}^{T}(m_{1}))$, $m$ also specifies the outcome for those who have matched in period $1$. Thus, part 4 should be read as “$(m_{t})_{t=2}^{T}$ coincides with an element of $\Sigma_{T-1}^{\Phi}(E^{T}_{2}(m_{1}))$ for the agents who are yet to be matched at the end of period 1.”
7 simply states that matching $m$ is a $\Phi$-solution for $E^{T}$ if the following two conditions hold. First, from period-2 onwards, $m$ specifies a $\Phi$-solution for the induced economy of length $T-1$, $E_{2}^{T}(m_{1})$. Second, $m$ has no period-1 blocks in the sense of 6. Because the correspondence $\Sigma_{T-1}^{\Phi}$ only depends on the conjectures from period $2$ onward, this recursive definition allows us in what follows to focus on the properties of $\varphi_{1}(\emptyset,\cdot)$, taking as given that the set $\Sigma_{T-1}^{\Phi}$ is nonempty.
@@ 4 Consistent conjectures
Given an economy $E^{T}$ and a matching $m\in\pazocal{M}_{T}$, the one-period economy induced by the collection $\{\varphi_{1}(m^{0},k):k\in A_{1}\cup B_{1}\}$ is the length-1 economy defined as follows:
@@
      For each agent $a\in A_{1}$, $u_{\varphi_{1}}(a,b)\geq u_{\varphi_{1}}(a,a)$ only if $u(a,b)\geq\min\{U_{1}(a,\overline{m}):\overline{m}\in\varphi_{1}(m^{0},a)\}$. Moreover, for $b,b^{\prime}\in B_{1}$, $u_{\varphi_{1}}(a,b)\geq u_{\varphi_{1}}(a,b^{\prime})$ only if $u(a,b)\geq u(a,b^{\prime})$.
@@
      For each agent $b\in B_{1}$, $v_{\varphi_{1}}(a,b)\geq v_{\varphi_{1}}(b,b)$ only if $v(a,b)\geq\min\{V_{1}(b,\overline{m}):\overline{m}\in\varphi_{1}(m^{0},b)\}$. Moreover, for $a,a^{\prime}\in A_{1}$, $v_{\varphi_{1}}(a,b)\geq v_{\varphi_{1}}(a^{\prime},b)$ only if $v(a,b)\geq v(a^{\prime},b)$.
@@
In words, a candidate matching, $m^{\star}\in M^{\star}$ satisfies the following. In period $1$, $m_{1}^{\star}$ is a stable matching for $E_{1,\varphi_{1}}$, i.e., $m^{\star}_{1}\in\mathrm{S}\left(E_{1,\varphi_{1}}\right)$. Because the one-period economy induced by $\varphi_{1}$ is devoid of any externalities, its set of stable matchings is nonempty and hence, at least one such matching exists (Gale and Shapley, 1962). Given the period-1 matching, $m^{\star}_{1}$, the continuation matching, $(m^{\star}_{t})_{t=2}^{T}$ is selected from $\Sigma_{T-1}^{\Phi}(E_{2}^{T}(m^{\star}_{1}))$, which is nonempty by assumption. Because $\mathrm{S}\left(E_{1,\varphi_{1}}\right)$ may not be a singleton and given one such stable matching, multiple continuation matchings may exist in the solution set of the induced economy, many candidates matchings may exist.Footnote 8Footnote 8Footnote 8Whereas many stable matchings for the one-period economy induced by the period-1 conjectures may exist, the Lone Wolf theorem implies they all have the same set of unmatched agents and hence, they all induce the same continuation economy.
Any such matching $m^{\star}$ is a natural candidate to be a $\Phi$-solution because it satisfies most of the conditions in 7. To see this, fix any such matching, $m^{\star}$, and note the following. First, $m^{\star}$ satisfies part 4 of 7 by construction. Second, $m^{\star}$ satisfies parts 1, 2, and 3 of 7 for all period-1 agents $k\in A_{1}\cup B_{1}$ such that $m^{\star}_{1}(k)\neq k$. By definition of $E_{1,\varphi_{1}}$, every agent who matches in $t=1$ is matched to someone better than their worst conjectured matching. Furthermore, because $m^{\star}_{1}$ is chosen from $\mathrm{S}\left(E_{1,\varphi_{1}}\right)$, no pair of matched agents under $m^{\star}_{1}$ can form a blocking pair.
@@
As we explain next, the only reason any such matching $m^{\star}$ may fail to be a $\Phi$-solution is that a “gap” exists between the conjectures of an agent who is unmatched in $t=1$ under $m^{\star}$ and the continuation matching $(m^{\star}_{t})_{t=2}^{T}$. To see this, note that the above discussion implies that if $m^{\star}$ is not a $\Phi$-solution, then $m^{\star}$ has a block involving a period-$1$ agent, $k$, who is unmatched in $t=1$, i.e., $m^{\star}_{1}(k)=k$. We claim next that if $k$ blocks $m^{\star}$, then $m^{\star}$ is not a conjecture for $k$. To fix ideas, let $k=a\in A_{1}$ and suppose $a$ blocks $m^{\star}_{1}$ with an agent $b\in B_{1}$. Because $m^{\star}_{1}\in\mathrm{S}\left(E_{1,\varphi_{1}}\right)$, it must be that at least one of $a$ or $b$ prefer their worst conjectured matching to matching with each other; assume for simplicity that this holds for $a$. Because $a$ wants to block $m^{\star}$ with $b$, it must be that $a$’s outcome under $m^{\star}$ is worse than the worst matching in $\varphi_{1}(\emptyset,a)$. In other words, when considering what matchings may ensue in the continuation economy, $a$ does not deem $m^{\star}$ a feasible conjecture. Because $a$ only cares about the continuation matching $(m^{\star}_{t})_{t=2}^{T}$, that $m^{\star}$ is not a feasible conjecture is akin to agent $a$’s conjectures not being consistent with the solution concept in the continuation. If instead $a$’s conjectures were consistent with the solution concept in the continuation, then they would not be willing to block $m^{\star}$ with $b$.
@@
Fix $E^{T}$ and a candidate matching $m^{\star}\in M^{\star}(\varphi_{1},\Sigma_{T-1}^{\Phi})$. Suppose that for all $k\in A_{1}\cup B_{1}$ such that $m^{\star}_{1}(k)=k$, we have that
@@
These conjectures reflect that $k$ believes that the $\Phi$-solutions from period $2$ onward describe the set of self-enforcing matchings for the induced economy. Thus, for each conjectured period-1 matching, $\overline{m}_{1}$, $k$ anticipates that the remaining unmatched agents and the new entrants will object to any matching $(\overline{m}_{t})_{t=2}^{T}$ that does not conform with $\Sigma_{T-1}^{\Phi}(\overline{m}_{1})$.Footnote 9Footnote 9Footnote 9These conjectures also appear in the solution concept for two-sided, repeated matching markets in Kotowski (2019). For instance, if $T=2$, these conjectures rule out matchings $\overline{m}$ such that the period-2 matching, $\overline{m}_{2}$ is not stable in the induced economy, $E_{2}^{2}(\overline{m}_{1})$. Interestingly, that these conjectures induce a nonempty solution concept implies that the perfection requirement alone is compatible with the existence of solutions for dynamic matching markets. Instead, difficulties may arise if the conjectures also place structure on the period-$t$ matching, as we do in Section 5.
@@
First, it is instructive to consider, in the context of the candidate matchings, the implications of not having consistent conjectures. To fix ideas, consider one such $m^{\star}$ and suppose an unmatched agent on side $A$, $a$, wishes to block $m^{\star}$ with an agent on side $B$, $b$. Suppose also that $a$ prefers the worst matching in $\varphi_{1}(\emptyset,a)$ to matching with $b$, and prefers matching with agent $b$ to $m^{\star}$, so that $m^{\star}\notin\varphi_{1}(\emptyset,a)$. There is a sense in which $a$’s block is not credible: Whereas $a$ prefers to match with $b$ over $m^{\star}$, as soon as $a$ matches with $b$, $a$ would prefer to break up with $b$. After all, $a$ prefers the worst matching in $\varphi_{1}(\emptyset,a)$ to matching with $b$. Thus, whenever the failure of consistency rules out a candidate matching $m^{\star}$ as a solution, this may be because the solution concept allows for these non-credible blocks.

### v2

@@ 1 Introduction
Formally, we study a two-sided, dynamic matching market, in which agents on both sides of the economy arrive over time. We assume arrivals are deterministic, that is, agents can perfectly foresee when each agent arrives in the economy.Footnote 2Footnote 2Footnote 2We assume deterministic arrivals because it simplifies notation, but it is not needed for our results. See 2.2. A matching specifies who matches with who in each period, subject to two constraints: agents cannot be matched before they arrive, and agents can have at most one matching partner on the opposite side. Agents are discounted utility maximizers, and their payoff from a matching solely depends on who they match with and when. Our goal is to study solution concepts for this economy.
@@
Whereas consistency is defined relative to the candidate matching, an immediate corollary is that the following conjectures always lead to a nonempty solution concept. Suppose that each agent’s conjectures only rule out continuation matchings that do not conform with the solution concept in the continuation. For instance, in a two-period economy, these conjectures specify that in the last period only stable matchings among the remaining unmatched agents and the new arrivals are possible. These conjectures always satisfy consistency and Theorem 4.1 implies they lead to a non-empty solution concept. That is, if the agents conjecture that the economy conforms to the solution concept in the continuation and place no restrictions on the current period matching, the set of matchings that conform to the solution concept from today onward is nonempty.
@@
Our paper builds on the literature on two-sided matching with externalities (Li, 1993; Sasaki and Toda, 1996; Hafalir, 2008) to contribute to the literature on two-sided dynamic matching markets with irreversible matching (Doval, 2022; Nicolò et al., 2023, 2024).
@@
The literature on cooperative solution concepts for dynamic matching markets highlights the role of agents’ conjectures about what matching would ensue after a block in defining solution concepts (see, e.g., Kurino, 2009; Kadam and Kotowski, 2018; Kotowski, 2019; Altinok, 2019; Liu, 2023 for repeated matching markets and Doval, 2022; Nicolò, Salmaso, and Saulle, 2023, 2024 for irreversible matching, and Doval (Forthcoming) for a survey treatment of this literature). Because solution concepts in this literature are typically predicated on perfection requirements–i.e., the same solution concept applying at the appropriately defined continuation–agents’ conjectures are defined jointly with the solution concept. We follow a complementary approach in which we take agents’ conjectures as given and use them to define a solution concept. This allows us to identify a property of conjectures–consistency–that guarantees the corresponding solution concept is nonempty, and underlies existence results in the literature (cf. Doval, 2022). We also propose new solution concepts that refine those existing in the literature, illustrating how our result can be used as a recipe of sorts to define stability notions for dynamic matching markets.
Recently, Nicolò, Salmaso, and Saulle (2024) study a third refinement of dynamic stability (Doval, 2022). Like the conjectures in dynamic stability, the conjectures in the solution concept in Nicolò et al. (2024) are defined recursively based on what the solution concept specifies in the continuation. In particular, the agents rule out all matchings that do not abide by the solution concept from tomorrow onward. Unlike dynamic stability, the solution concept in Nicolò et al. (2024) imposes more structure on what matchings the agents can conjecture will ensue in the current period when considering remaining unmatched: The authors define conjectures in a given period through an iterative procedure which, at each stage, rules out current period matchings that do not satisfy some minimal stability requirements given the conjectures at that stage.Footnote 3Footnote 3Footnote 3Whereas the main focus of the paper is on the conjectures that obtain as the limit of this procedure, by stopping the procedure at a given stage the authors can study stability under different levels of sophistication (Hafalir, 2008). It is possible to show the conjectures they construct satisfy consistency. Following Doval (2022), their proof is based on showing that (the analogue for their solution concept of) the candidate matchings we define in Section 4 satisfy the conditions of their solution concept. Furthermore, the authors show via an example that the rational expectations solution may fail to exist in a two-sided dynamic matching market.
@@ 2 Model
A length-$T$ economy, denoted $E^{T}=(A_{1},B_{1},\dots,A_{T},B_{T})$, consists of two sequences, $\{A_{1},\dots,A_{T}\}$ and $\{B_{1},\dots,B_{T}\}$, of disjoint subsets of $\mathscr{A}$ and $\mathscr{B}$, respectively.Footnote 4Footnote 4Footnote 4It is without loss to assume the sets $(A_{1},\dots,A_{T})$ are disjoint as we do not rule out indifferences. Thus, $a$ denotes a unique individual in the economy, but $a$ may be payoff-equivalent to other arriving agents on side $A$. The same applies for the sets $(B_{1},\dots,B_{T})$ being disjoint. For $t\in\{1,\dots,T\}$, $A_{t}$ are the side $A$-agents that arrive in period $t$, and $B_{t}$ are the side $B$-agents that arrive in period $t$. For any $t\leq T$, let $\overline{A}_{t}=\cup_{s=1}^{t}A_{s}$ denote the implied arrivals on side $A$ through period $t$; similarly, let $\overline{B}_{t}=\cup_{s=1}^{t}B_{s}$ denote the implied arrivals on side $B$ through period $t$.
@@
We close the model by defining the agents’ preferences. Throughout, we assume agents discount the future. Formally, each $a\in\mathscr{A}$ defines a discount factor $\delta_{a}\in[0,1]$ and a Bernoulli utility, $u(a,\cdot):\mathscr{B}\cup\{a\}\mapsto\mathbb{R}$. Similarly, each $b\in\mathscr{B}$ defines a discount factor $\delta_{b}\in[0,1]$ and a Bernoulli utility, $v(\cdot,b):\mathscr{A}\cup\{b\}\mapsto\mathbb{R}$. We normalize to 00 the payoffs of remaining single for one period. That is, for all $a\in\mathscr{A}$ and all $b\in\mathscr{B}$, $u(a,a)=v(b,b)=0$.
@@
Remark 2.2 (Simplifying assumptions).
We note two assumptions that simplify notation, but are otherwise not needed for the results. First, we assume agents can perfectly foresee who arrives in each period. Instead, our results extend to the case in which arrivals are stochastic so that an economy is defined as a distribution over sequences $E^{T}$ as in Doval (2021). Second, we assume time-discounted preferences, but all that matters for our results is that agents are dynamically consistent.
@@ 3 A solution concept given conjectures
Because the conjectures represent what $k$ expects to obtain should they remain unmatched in period $t$, each conjectured matching $\overline{m}$ satisfies $\overline{m}_{t}(k)=k$. Each conjectured matching $\overline{m}\in\varphi_{t}(m^{t-1},k)$ describes both the matching that ensues in period $t$, $\overline{m}_{t}$, and the continuation matching $(\overline{m}_{t+1},\dots,\overline{m}_{T})$. Whereas agent $k$’s matching outcome is determined by the continuation matching $(\overline{m}_{t+1},\dots,\overline{m}_{T})$, specifying who matches in period $t$ is necessary to define what matchings are feasible from period $t+1$ onward. For this reason, a conjecture describes the matching starting from period $t$ onward. That defining the set of matchings $k$ may face from period $t+1$ onward necessitates specifying the outcomes of the other period-$t$ agents implies an externality exists in this economy, even if no direct payoff externalities exist.
@@
In words, a matching $m$ is a $\Phi$-solution for $E^{T}$ if in each period the following holds. First, each agent $k$ who is available to match in period $t$ prefers $m$ to their worst conjectured matching (parts 1 and 2 in 6). Note that this implies that if $m_{t}(k)\neq k$, then $k$ prefers their matching partner to waiting to be matched, and if $m_{t}(k)=k$, then $k$ has no objections to the continuation matching, $(m_{s})_{s=t+1}^{T}$.Footnote 5Footnote 5Footnote 5In a static economy, only the first kind of block may exist. The second block only shows up in a dynamic economy (see Doval, 2022 for a discussion). For instance, if the continuation matching is not self-enforcing, an agent may object to it. As we discuss later on, consistency of the conjectures rules is very much related to the second class of blocks. In both cases, we say $k$ prefers to be available to match in period $t$. Second, no pair of agents who can match in period $t$ prefer to match with each other to matching according to $m$ (part 3 in 6). Note that $(a,b)$ being able to match in period $t$ does not imply $a$ and $b$ arrive in the same period, but that under matching $m$ they meet in period $t$ and are able to block $m$.
@@
Whereas 6 is fairly general, it is worth noting that the definition rules out solution concepts in which the conjectures depend on the matching under consideration beyond the induced economy in period $t$. For instance, when considering whether to block matching $m$ in period $t$, if agent $k$ presumes all other contemporaneous agents–except perhaps for $k$’s matching partner–remain matched according to $m_{t}$, then $k$’s conjectures depend on the outcome specified by $m$ in period $t$. (These conjectures are referred to as passive conjectures; see Chowdhury (2004) and Doval (2015).) Instead, 6 presumes we can define the conjectures without reference to any specific matching. The matching under consideration only matters inasmuch as it determines who can match in each period, and hence whose conjectures one should check in each period.
@@
    For all $a\in A_{1}$, a matching $\overline{m}\in\varphi_{1}(\emptyset,a)$ exists such that $U_{1}(a,m)\geq U_{1}(a,\overline{m})$,
@@
    For all $b\in B_{1}$, a matching $\overline{m}\in\varphi_{1}(\emptyset,b)$ exists such that $V_{1}(b,m)\geq V_{1}(b,\overline{m})$,
@@
    $(m_{t})|_{t=2}^{T}\in\Sigma_{T-1}^{\Phi}(E_{2}^{T}(m_{1}))$.Footnote 7Footnote 7Footnote 7We abuse notation slightly. Whereas $\Sigma_{T-1}^{\Phi}(E_{2}^{T}(m_{1}))$ defines a matching only for the agents who can match from $t=2$ onward–i.e., it is an element of $\pazocal{M}_{T-1}(E_{2}^{T}(m_{1}))$$m$ also specifies the outcome for those who have matched in period $1$. Thus, part 4 should be read as “$(m_{t})_{t=2}^{T}$ coincides with an element of $\Sigma_{T-1}^{\Phi}(E^{T}_{2}(m_{1}))$ for the agents who are yet to be matched at the end of period 1, and coincides with $m_{1}$, otherwise.”
7 simply states that matching $m$ is a $\Phi$-solution for $E^{T}$ if the following two conditions hold. First, from period $2$ onwards, $m$ specifies a $\Phi$-solution for the induced economy of length $T-1$, $E_{2}^{T}(m_{1})$. Second, $m$ has no blocks in period $1$ in the sense of 6. Because the correspondence $\Sigma_{T-1}^{\Phi}$ only depends on the conjectures from period $2$ onward, this recursive definition allows us in what follows to focus on the properties of $\varphi_{1}(\emptyset,\cdot)$, taking as given that the set $\Sigma_{T-1}^{\Phi}$ is nonempty.
@@ 4 Consistent conjectures
Given an economy $E^{T}$ and a matching $m\in\pazocal{M}_{T}$, the one-period economy induced by the collection $\{\varphi_{1}(\emptyset,k):k\in A_{1}\cup B_{1}\}$ is the length-1 economy defined as follows:
@@
      For each agent $a\in A_{1}$, $u_{\varphi_{1}}(a,b)\geq u_{\varphi_{1}}(a,a)$ only if $u(a,b)\geq\min\{U_{1}(a,\overline{m}):\overline{m}\in\varphi_{1}(\emptyset,a)\}$. Moreover, for $b,b^{\prime}\in B_{1}$, $u_{\varphi_{1}}(a,b)\geq u_{\varphi_{1}}(a,b^{\prime})$ only if $u(a,b)\geq u(a,b^{\prime})$.
@@
      For each agent $b\in B_{1}$, $v_{\varphi_{1}}(a,b)\geq v_{\varphi_{1}}(b,b)$ only if $v(a,b)\geq\min\{V_{1}(b,\overline{m}):\overline{m}\in\varphi_{1}(\emptyset,b)\}$. Moreover, for $a,a^{\prime}\in A_{1}$, $v_{\varphi_{1}}(a,b)\geq v_{\varphi_{1}}(a^{\prime},b)$ only if $v(a,b)\geq v(a^{\prime},b)$.
@@
In words, a candidate matching, $m^{\star}\in M^{\star}$ satisfies the following. In period $1$, $m_{1}^{\star}$ is a stable matching for $E_{1,\varphi_{1}}$, i.e., $m^{\star}_{1}\in\mathrm{S}\left(E_{1,\varphi_{1}}\right)$. Because the one-period economy induced by $\varphi_{1}$ is devoid of any externalities, its set of stable matchings is nonempty and hence, at least one such matching exists (Gale and Shapley, 1962). Given the period-1 matching, $m^{\star}_{1}$, the continuation matching, $(m^{\star}_{t})_{t=2}^{T}$ is selected from $\Sigma_{T-1}^{\Phi}(E_{2}^{T}(m^{\star}_{1}))$, which is nonempty by assumption. Because $\mathrm{S}\left(E_{1,\varphi_{1}}\right)$ may not be a singleton and given one such stable matching, multiple continuation matchings may exist in the solution set of the induced economy, many candidates matchings may exist.Footnote 8Footnote 8Footnote 8Whereas many stable matchings for the one-period economy induced by the period-1 conjectures may exist, the Lone Wolf theorem (McVitie and Wilson, 1970) implies they all have the same set of unmatched agents and hence, they all induce the same continuation economy.
Any matching in $M^{\star}(\varphi_{1},\Sigma_{T-1}^{\Phi})$ is a natural candidate to be a $\Phi$-solution because it satisfies most of the conditions in 7. To see this, fix any such matching, $m^{\star}$, and note the following. First, $m^{\star}$ satisfies part 4 of 7 by construction. Second, $m^{\star}$ satisfies parts 1, 2, and 3 of 7 for all period-1 agents $k\in A_{1}\cup B_{1}$ such that $m^{\star}_{1}(k)\neq k$. By definition of $E_{1,\varphi_{1}}$, every agent who matches in $t=1$ is matched to someone better than their worst conjectured matching. Furthermore, because $m^{\star}_{1}$ is chosen from $\mathrm{S}\left(E_{1,\varphi_{1}}\right)$, no pair of matched agents under $m^{\star}_{1}$ can form a blocking pair.
@@
As we explain next, the only reason any such matching $m^{\star}$ may fail to be a $\Phi$-solution is that a “gap” exists between the conjectures of an agent who is unmatched in $t=1$ under $m^{\star}$ and the matching $m^{\star}$. To see this, note that the above discussion implies that if $m^{\star}$ is not a $\Phi$-solution, then $m^{\star}$ has a block involving a period-$1$ agent, $k$, who is unmatched in $t=1$, i.e., $m^{\star}_{1}(k)=k$. We claim next that if $k$ blocks $m^{\star}$, then $m^{\star}$ is not a conjecture for $k$. To fix ideas, let $k=a\in A_{1}$ and suppose $a$ blocks $m^{\star}_{1}$ with an agent $b\in B_{1}$. Because $m^{\star}_{1}\in\mathrm{S}\left(E_{1,\varphi_{1}}\right)$, it must be that at least one of $a$ or $b$ prefer their worst conjectured matching to matching with each other; assume for simplicity that this holds for $a$. Because $a$ wants to block $m^{\star}$ with $b$, it must be that $a$’s outcome under $m^{\star}$ is worse than the worst matching in $\varphi_{1}(\emptyset,a)$. In other words, when considering what matchings may ensue in the continuation economy, $a$ does not deem $m^{\star}$ a feasible conjecture. Because $a$ only cares about the continuation matching $(m^{\star}_{t})_{t=2}^{T}$, that $m^{\star}$ is not a feasible conjecture is akin to agent $a$’s conjectures not being consistent with the solution concept in the continuation. If instead $a$’s conjectures were consistent with the solution concept in the continuation, then they would not be willing to block $m^{\star}$ with $b$.
@@
Suppose $\Sigma_{T-1}^{\Phi}$ is a nonempty-valued correspondence. Fix $E^{T}$ and a candidate matching $m^{\star}\in M^{\star}(\varphi_{1},\Sigma_{T-1}^{\Phi})$. Suppose that for all $k\in A_{1}\cup B_{1}$ such that $m^{\star}_{1}(k)=k$, we have that
@@
These conjectures reflect that $k$ believes that the economy abides by the set of $\Phi$-solutions from period $2$ onward. Thus, for each conjectured period-1 matching, $\overline{m}_{1}$, $k$ anticipates that the remaining unmatched agents and the new entrants will object to any matching $(\overline{m}_{t})_{t=2}^{T}$ that does not conform with $\Sigma_{T-1}^{\Phi}(\overline{m}_{1})$.Footnote 9Footnote 9Footnote 9These conjectures also appear in the solution concept for two-sided, repeated matching markets in Kotowski (2019). For instance, if $T=2$, these conjectures rule out matchings $\overline{m}$ such that the period-2 matching, $\overline{m}_{2}$ is not stable in the induced economy, $E_{2}^{2}(\overline{m}_{1})$. Interestingly, that these conjectures induce a nonempty solution concept implies that the perfection requirement alone is compatible with the existence of solutions for dynamic matching markets. Instead, difficulties may arise if the conjectures also place structure on the period-$t$ matching, as we do in Section 5.
@@
First, it is instructive to consider the implications of not having consistent conjectures in the context of the candidate matchings. To fix ideas, consider one such $m^{\star}$ and suppose an unmatched agent on side $A$, $a$, wishes to block $m^{\star}$ with an agent on side $B$, $b$. Suppose also that $a$ prefers the worst matching in $\varphi_{1}(\emptyset,a)$ to matching with $b$, and prefers matching with agent $b$ to $m^{\star}$, so that $m^{\star}\notin\varphi_{1}(\emptyset,a)$. There is a sense in which $a$’s block is not credible: Whereas $a$ prefers to match with $b$ over $m^{\star}$, as soon as $a$ matches with $b$, $a$ would prefer to break up with $b$. After all, $a$ prefers the worst matching in $\varphi_{1}(\emptyset,a)$ to matching with $b$. Thus, whenever the failure of consistency rules out a candidate matching $m^{\star}$ as a solution, this may be because the solution concept allows for these non-credible blocks.